# Value Flows

### v1

@@ C.1 Practical flow matching losses
Our practical loss for fitting the return distribution involves the DCFM loss and a bootstrapped regularization based on TD learning. Since it is computationally inefficient to keep track of a single historical vector field $v_{k}$ and optimize the new vector field $v$ until convergence, we will use a target vector field $\bar{v}$ (Farebrother et al., 2025), which aggregates all the historical vector fields $\{v_{k}\}$, to replace the single historical vector field in ${\mathcal{L}}_{\text{DCFM}}$ (Eq. 15). Using $z^{t}\sim\bar{p}(z^{t}\mid t,s^{\prime},a^{\prime})$ to denote the sampling procedure of (1) first sampling a noise $\epsilon\sim{\mathcal{N}}(0,1)$ and (2) then invoking the Euler method (Alg. 2), we have
${\mathcal{L}}_{\text{DCFM}}(v)=\mathbb{E}_{\begin{subarray}{c}(s,a,r,s^{\prime})\sim D,t\sim\textsc{Unif}([0,1])\\ a^{\prime}\sim\pi(a^{\prime}\mid s^{\prime}),z^{t}\sim\bar{p}(z^{t}\mid t,s^{\prime},a^{\prime})\end{subarray}}\left[\left(v(r+\gamma z^{t}\mid t,s,a)-v_{k}\left(z^{t}\mid t,s^{\prime},a^{\prime}\right)\right)^{2}\right].$ (17)
However, in our initial experiments, naively optimizing ${\mathcal{L}}_{\text{DCFM}}(v)$ produced a divergent vector field. To stabilize learning, we use the return predictions at the next state-action pair $(s^{\prime},a^{\prime})$ and the flow time $t=1$ to construct a bootstrapped target return and invoke the standard conditional flow matching loss (Eq. 13). The resulting loss function, called bootstrapped conditional flow matching (BCFM) loss, resembles the standard Bellman error. Specifically, using $z^{1}_{\text{TD}}=r(s,a)+\gamma z^{1}$ to denote the target return, and using $z^{t}_{\text{TD}}=tz^{1}_{\text{TD}}+(1-t)\epsilon$ to denote a linear interpolation between $z^{1}$ and $\epsilon$, [^5] The same noise $\epsilon$ is used to sample $z^{1}$ and construct $z^{t}_{\text{TD}}$ and $z^{1}_{\text{TD}}-\epsilon$. the BCFM loss can be written as
@@ C.2 Policy extraction strategies
Second, for online fine-tuning in offline-to-online RL, following prior work (Park et al., 2025c), we learn a stochastic one-step policy $\pi:{\mathcal{S}}\times\mathbb{R}^{d}\to{\mathcal{A}}$ to minimize a DDPG-style loss with behavioral regularization (Fujimoto & Gu, 2021). This loss function guides the policy to select actions that maximize the Q estimates while distilling it toward the fixed BC flow policy:
@@
Appendix D Full Experiment Results
D.1 Offline Results
@@
relocate-expert-v1 $\mathbf{108}$ $\mathbf{106}$ $\mathbf{108}$ $\mathbf{105}\pm 2$ $\mathbf{104}\pm 3$ $\mathbf{107}\pm 1$ $\mathbf{103}\pm 0$ $\mathbf{103}\pm 0$ $\mathbf{103}\pm 2$ $102\pm 2$
@@ D.2 Offline-to-Online Results
D.3 Analysis on ablations
Figure 7: Value Flows is robust against the number of flow steps in the Euler method.
@@ D.3 Analysis on ablations
In this section, we conduct full ablation experiments to study the effects of (1) the regularization loss coefficient, (2) the confidence weight, (3) the number of flow steps in the Euler method, and (4) the number of candidates in rejection sampling.
To better understand the role of the BCFM regularization loss, we ablate over different values of the regularization coefficient $\lambda$ and compare the performance of Value Flows. We present the results in Fig. 5. We observe that removing the BCFM regularization loss ($\lambda=0$) results in poor performance on both cube-double and puzzle-3x3, while Value Flows’s performance starts to increase after adding some regularizations. Choosing the correct regularization coefficient $\lambda$ boosts the mean performance of Value Flows by $2.6\times$ and increases the sample efficiency of our algorithm by $2\times$. This suggests the BCFM regularization is an important component of Value Flows.
A key design choice of Value Flows is the confidence weight. We next discuss the effect of using different temperatures $\tau$ in the confidence weight. As shown in Fig. 5, we observe that, on puzzle-3x3 tasks, using a constant confidence weight ($\tau=0$) can result in slower convergence of the success rate. On cube-double tasks, using a constant confidence weight results in slightly worse performance. We see that an appropriate temperature $\tau$ boosts the success rate by $60\%$ on average. Comparing against different magnitudes of the confidence weighting, we find that our choice in Table 4 is optimal.
@@
Our final experiments study the effects of the number of flow steps in the Euler method and the number of candidates in rejection sampling. We ablate over these hyperparameters and present the results in Fig. 8. Note that we keep the number of flow steps for the flow policy fixed, and only vary the number of flow steps in the return vector field. We observe that, on puzzle-3x3 tasks, using a small number of flow steps can drastically decrease the success rate. Specifically, using only $5$ flow steps resulted in a $20\%$ decrease in success rate compared to the results with $10$ flow steps. We conjecture that the number of flow steps affects the expressiveness and accuracy of the return distribution. Empirically, we found that $10$ flow steps worked well and kept it fixed throughout our experiments.
We also study the effect of the number of candidates in rejection sampling for policy extraction. Results in Fig. 8 show that Value Flows is robust to the number of candidates in rejection samples, except that using $32$ candicates results in $30\%$ decrease in success rate on puzzle-3x3 tasks. In our experiments, we found that $16$ candidates are sufficient across a variety of tasks.
Appendix E Visualizing the confidence weight
Appendix F Experiment details
@@
F.1 Environments
We use OGBench (Park et al., 2025a) and D4RL (Fu et al., 2021) benchmarks to perform experimental evaluations. OGBench is a benchmark for offline goal-conditioned reinforcement learning, with single-task variants for standard reward-maximizing RL algorithms. We choose the single-task variants for all domains in Table 2. For locomotion tasks, the reward is either 1-1 or 0, where 0 indicates task completion. For manipulation tasks, the reward varies between $n_{\text{task}}$ and 0, depending on how many subtasks are complete. We use the following datasets from OGBench for each domain.
@@ F.1 Environments
D4RL is a popular benchmark for studying offline RL. We measure the performance by the normalized returns following Fu et al. (2021). We use the following 12 tasks from Adroit, which require dexterous manipulation within a 24-DoF action space.
@@
The cube-double and cube-triple tasks involve complex pick-and-place of colored cube blocks. The scene tasks involve long-horizon reasoning and interaction with various objects in the scene. The puzzle-3x3 and puzzle-4x4 tasks require solving the "Lights Out" puzzles using the robot arm and further test combinatorial generalization. The Adroit benchmarks involve controlling a 28-DoF hand to perform dexterous skills: spinning a pen, opening a door, relocating a ball, and using a hammer to knock a button. The antmaze and humanoidmaze tasks are designed to navigate a challenging maze by controlling an 8-DoF ant and a 21-DoF humanoid respectively. The visual variants of these navigation tasks further involve working with a partially observable MDP, as there is no low-dimensional state information, and the algorithm must directly learn from $64\times 64\times 3$ RGB images. We include challenging state-based and image-based variants for manipulation tasks as they require tackling multimodal returns and long-horizon reasoning. We also include some image-based navigation tasks to evaluate the algorithms under partially observable settings.
F.2 Methods for Comparison
We compare Value Flows against 9 baselines, measuring the performance using success rates. Specifically, we first compare to BC (Fujimoto & Gu, 2021), IQL (Kostrikov et al., 2021a), and ReBRAC (Tarasov et al., 2023) as representative methods that learn scalar Q values with a Gaussian policy. The second set of baselines is state-of-the-art methods that learn scalar Q values with a flow policy (FBRAC, IFQL, FQL) (Park et al., 2025c). We also include comparisons against prior distributional RL methods that model the return distribution as a categorical distribution (C51 (Bellemare et al., 2017)) or a finite number of quantiles (IQN (Dabney et al., 2018), and CODAC (Ma et al., 2021)). We use a flow policy for all these distributional RL baselines. Following prior work (Park et al., 2025c), we report means and standard deviations over 8 random seeds for state-based tasks (4 seeds for image-based tasks). We describe the details of each baseline below.

### v2

@@ C.1 Practical flow matching losses
Our practical loss for fitting the return distribution involves the DCFM loss and a bootstrapped regularization based on TD learning.
We will use a target vector field $\bar{v}$ (Farebrother et al., 2025b), which aggregates all the historical vector fields $\{v_{k}\}$, to replace the single historical vector field in ${\mathcal{L}}_{\text{DCFM}}$ (Eq. 15). The reasons for introducing the target vector field $\hat{v}$ are twofold. First, using a target network is a widely used technique in RL. This technique has been proven to smooth learning and enable standard TD learning (Mnih et al., 2015; Farebrother et al., 2025a). Second, a vector field minimizing the preliminary loss ${\mathcal{L}}_{\text{DCFM}}$ might collapse to a constant, e.g., $v=0$, because the vector field is fitting the output of itself. However, the same problem has already existed in a family of self-supervised representation learning methods, e.g., BYOL (Grill et al., 2020), where they learn representations to predict output from the model itself. Instead of using the model from the previous iteration, they use an exponential moving average of the historical models (target model) to predict target outputs. In the experiments of Grill et al. (2020), this target network helps prevent model collapse, and this observation motivates us to use the target vector field in Value Flows. Additionally, Tian et al. (2021) provides a theoretical characterization of how BYOL avoids this failure mode. Using $z^{t}\sim\bar{p}(z^{t}\mid t,s^{\prime},a^{\prime})$ to denote the sampling procedure of (1) first sampling a noise $\epsilon\sim{\mathcal{N}}(0,1)$ and (2) then invoking the Euler method (Alg. 2), we have
${\mathcal{L}}_{\text{DCFM}}(v)=\mathbb{E}_{\begin{subarray}{c}(s,a,r,s^{\prime})\sim D,t\sim\textsc{Unif}([0,1])\\ a^{\prime}\sim\pi(a^{\prime}\mid s^{\prime}),z^{t}\sim\bar{p}(z^{t}\mid t,s^{\prime},a^{\prime})\end{subarray}}\left[\left(v(r+\gamma z^{t}\mid t,s,a)-\bar{v}\left(z^{t}\mid t,s^{\prime},a^{\prime}\right)\right)^{2}\right].$ (17)
In our initial experiments, naively optimizing ${\mathcal{L}}_{\text{DCFM}}(v)$ produced a trivial performance on some tasks. To stabilize learning, we use the return predictions at the next state-action pair $(s^{\prime},a^{\prime})$ and the flow time $t=1$ to construct a bootstrapped target return and invoke the standard conditional flow matching loss (Eq. 13). The resulting loss function, called bootstrapped conditional flow matching (BCFM) loss, resembles the standard Bellman error. Specifically, using $z^{1}_{\text{TD}}=r(s,a)+\gamma z^{1}$ to denote the target return, and using $z^{t}_{\text{TD}}=tz^{1}_{\text{TD}}+(1-t)\epsilon$ to denote a linear interpolation between $z^{1}$ and $\epsilon$, [^4] The same noise $\epsilon$ is used to sample $z^{1}$ and construct $z^{t}_{\text{TD}}$ and $z^{1}_{\text{TD}}-\epsilon$. the BCFM loss can be written as
@@ C.2 Policy extraction strategies
Second, for online fine-tuning in offline-to-online RL, following prior work (Park et al., 2025c), we learn a flow one-step policy $\pi:{\mathcal{S}}\times\mathbb{R}^{d}\to{\mathcal{A}}$ to minimize a DDPG-style loss with behavioral regularization (Fujimoto & Gu, 2021). This loss function guides the policy to select actions that maximize the Q estimates while distilling it toward the fixed BC flow policy:
@@
Appendix D Additional experiments
@@
relocate-expert-v1 $\mathbf{108}$ $\mathbf{106}$ $\mathbf{108}$ $\mathbf{105}\pm 2$ $\mathbf{104}\pm 3$ $\mathbf{107}\pm 1$ $\mathbf{103}\pm 0$ $\mathbf{103}\pm 0$ $\mathbf{103}\pm 2$ $\mathbf{105}\pm 3$
@@
D.1 Analysis on ablations
Figure 7: Value Flows is robust against the number of flow steps in the Euler method for the return vector field after a certain threshold ($10$).
@@ D.1 Analysis on ablations
In this section, we conduct full ablation experiments to study the effects of (1) the regularization loss coefficient, (2) the confidence weight, (3) the number of flow steps in the Euler method, and (4) the number of candidates in rejection sampling. Again, we compute means and standard deviations over $8$ random seeds.
To better understand the role of the BCFM regularization loss, we ablate over different values of the regularization coefficient $\lambda$ and compare the performance of Value Flows. We present the results in Fig. 5. We observe that removing the BCFM regularization loss ($\lambda=0$) results in poor performance on both cube-double-play-singletask-task2-v0 and puzzle-3x3-play-singletask-task4-v0, while Value Flows’s performance starts to increase after adding some regularizations. Choosing the correct regularization coefficient $\lambda$ boosts the mean performance of Value Flows by $2.6\times$ and increases the sample efficiency of our algorithm by $2\times$. This suggests the BCFM regularization is an important component of Value Flows.
A key design choice of Value Flows is the confidence weight. We next discuss the effect of using different temperatures $\tau$ in the confidence weight. As shown in Fig. 5, we observe that, on the puzzle-3x3-play-singletask-task4-v0 task, using a constant confidence weight ($\tau=0$) can result in slower convergence of the success rate. On the cube-double-play-singletask-task2-v0 task, using a constant confidence weight results in slightly worse performance. We see that an appropriate temperature $\tau$ boosts the success rate by $60\%$ on average. Comparing against different magnitudes of the confidence weighting, we find that our choice in Table 6 is optimal.
Our final experiments study the effects of the number of flow steps in the Euler method and the number of candidates in rejection sampling. We ablate the number of flow steps within $\{1,2,5,10,20,30\}$ and ablate the number of rejection sampling candidates within $\{8,16,32,64\}$, presenting the results in Fig. 8. Note that we keep the number of flow steps for the flow policy fixed, and only vary the number of flow steps when sampling from the learned return model. We observe that, on the cube-double-play-singletask-task2-v0 task, even using a very small number of flow steps ($1$ or $2$) does not significantly decrease the performance of Value Flows. In contrast, using a small number of flow steps on the puzzle-3x3-play-singletask-task4-v0 task can drastically reduce the success rate. Specifically, using $5$ flow steps resulted in a $20\%$ decrease in success rate compared to the results with $10$ flow steps, while using $1$ or $2$ flow steps will result in near-zero success rates. We also observe that the performance saturates after increasing the number of flow steps to $10$. We conjecture that the number of flow steps affects the expressiveness and accuracy of the return distribution. Empirically, we found that $10$ flow steps worked well and kept it fixed throughout our experiments.
We also study the effect of the number of candidates in rejection sampling for policy extraction. Results in Fig. 8 show that Value Flows is robust to the number of candidates in rejection samples, except that using $32$ candicates results in a $30\%$ decrease in success rate on the puzzle-3x3-play-singletask-task4-v0 task. In our experiments, we found that $16$ candidates are sufficient across a variety of tasks.
D.2 Visualizing the confidence weight
D.3 Computational time and resource consumption
@@
Figure 10: Computational time and resource usage of different algorithms. Training Value Flows requires intermediate wall time and GPU memory compared with prior methods.
We analyzed our experiment logs to investigate further the computational time and GPU memory consumption of different methods. Specifically, we analyze experiment logs on one state-based OGBench task (puzzle-4x4-play-singletask-task4-v0) and one image-based OGBench task (visual-cube-double-play-singletask-task1-v0), comparing Value Flows to selection baselines: IQL, ReBRAC, FBRAC, IFQL, FQL, and IQN. We report wall-time and maximum GPU memory usage aggregated across 8 seeds (4 seeds for the image-based task) for different methods. Results in Fig. 10 suggest that training Value Flows requires a computational time similar to training FQL, and the mean wall time of Value Flows is $50\%$ of the mean wall time of IQN. To maximize GPU usage, Value Flows used substantially less GPU memory than FBRAC and IQN because they directly use the Euler method in the actor loss (FBRAC) or employ additional neural networks to encode quantiles (IQN). Taken together, training Value Flows requires intermediate wall time and GPU memory compared with prior methods.
D.4 Solving a risk-sensitive MDP
$s_{0}$$s_{1}$$s_{9}$$s_{10}$${\mathcal{N}}(0,10^{-4})$${\mathcal{N}}(0,10^{-4})$${\mathcal{N}}(-130,20)$${\mathcal{N}}(-130,20)$${\mathcal{N}}(-130,20)$${\mathcal{N}}(-100,800)$replacenot replaceterminal state
Figure 11: The machine replacement MDP. (Left)  The machine replacement MDP consists of two actions replace and not replace with the reward distributed as Gaussians. This MDP is challenging because it requires exploration and risk-sensitive control. (Right)Value Flows converges to the optimal $\text{CVaR}_{0.1}$ with a number of gradient steps similar to IQN, while C51 failed to converge within $5000$ gradient steps. The Q-learning immediately converges to the optimal expected return and is not able to learn a risk-sensitive policy.
We now use a domain that requires exploration and risk-sensitive control to demonstrate the capability of Value Flows. Following prior work (Keramati et al., 2020), we choose a variant of the Machine Replacement MDP (with only $11$ states) to conduct our experiments (Fig. 11 (Left)), measuring the $\text{CVaR}_{0.1}$ performance instead of the expected return. This MDP is especially challenging because of the chain structure, and it also requires risk-sensitive control due to the high variance in the reward distributions at the end of the chain. Therefore, it serves as a good evaluation task for distributional RL algorithms. Note that the optimal policy for expected return (Q-optimal policy) always chooses not replace, taking the risk of high variance reward at $s_{9}$. In contrast, the optimal policy for maximizing $\text{CVaR}_{0.1}$ ($\text{CVaR}_{0.1}$-optimal policy) will choose to replace at state $s_{9}$ to avoid the high variance reward.
Since the machine replacement MDP has a discrete action space, we remove the policy extraction procedure in Value Flows and use the learned vector fields to choose greedy actions directly, resembling the original C51 (Bellemare et al., 2017) and IQN (Dabney et al., 2018) implementations. We use a random policy and uniformly initialize the start state to collect a dataset with $10^{5}$ transitions for offline training. We compare Value Flows against C51 and IQN and use $\text{CVaR}_{0.1}$ instead of the Q-value ($\text{CVaR}_{0.5}$) to optimize all methods. We compute the means and standard deviations of $\text{CVaR}_{0.1}$ over 4 random seeds. Results in Fig. 11 (Right) show that Value Flows converges to the optimal $\text{CVaR}_{0.1}$ with the number of gradient steps similar to IQN ($600$), while C51 failed to converge within $5000$ gradient steps. We also observe that Q-learning immediately converges to the optimal expected return and is not able to learn a risk-sensitive policy.
D.5 Using Gaussian policies on D4RL adroit tasks
Figure 12: Using a Gaussian BC policy instead of a flow BC policy for rejection sampling improves the performance of Value Flows on D4RL adroit tasks by $10\%$, matching the state-of-the-art performance achieved by ReBRAC.
Figure 13: Training Value Flows only using the BCFM loss yields near-zero success rates. In contrast, minimizing the DCFM loss provides the most learning signal for Value Flows.
We hypothesize that flow-based policies generally performed worse than Gaussian policies on D4RL adroit tasks (see Table 1 and Table 3). This hypothesis is consistent with the experiment results in prior work, e.g., Table 2 of Park et al. (2025c). The D4RL adroit tasks consist of a very narrow dataset distribution: small amounts of human demonstrations, large amounts of expert data from a fine-tuned RL policy, and a mixture of human demonstrations and behavioral cloning trajectories (Fu et al., 2021). We conjecture that this narrow dataset distribution might impose difficulty for the expressive flow-matching policy to learn a multi-modal action distribution, resulting in sampling out-of-distribution actions during policy extraction.
To test this hypothesis, we ran additional ablations to study the effects of using Gaussian policies in Value Flows on D4RL adroit tasks. Instead of learning a flow BC policy as in Appendix C.2, we learn a Gaussian BC policy and then apply rejection sampling. We choose to conduct experiments on the four expert variants of D4RL adroit tasks: pen-expert-v1, door-expert-v1, hammer-expert-v1, and relocate-expert-v1, aggregating results over 8 seeds. Results in Fig. 13 suggest that using Gaussian BC policies improves the mean performance of Value Flows by $10\%$, matching the state-of-the-art performance achieved by ReBRAC.
D.6 Role of the bootstrapped conditional flow matching loss
In the complete Value Flows flow matching objective, we include a BCFM regularization loss ${\mathcal{L}}_{\text{wBCFM}}$ to stabilize learning. This section studies whether the DCFM loss ${\mathcal{L}}_{\text{wDCFM}}$ or the BCFM loss ${\mathcal{L}}_{\text{wBCFM}}$ provides the most learning signal for Value Flows. We hypothesize that simply using the BCFM loss to train the vector field in Value Flows does not provide a sufficient learning signal to model the return distribution.
To test this hypothesis, we conduct ablation experiments comparing a variant of Value Flows trained only using the BCFM loss against the original Value Flows. We evaluate the success rates on two OGBench tasks cube-triple-play-singletask-task1-v0 and scene-play-singletask-task2-v0, computing means and standard deviations of success rates after training over $8$ seeds. Results in Fig. 13 suggest that training Value Flows only using the BCFM loss yields near-zero success rates on both tasks. In contrast, minimizing the DCFM loss provides the most learning signal for Value Flows. Thus, the BCFM loss mainly serves as a regularization.
Appendix E Experiment details
@@
E.1 Environments
We use OGBench (Park et al., 2025a) and D4RL (Fu et al., 2021) benchmarks to perform experimental evaluations. OGBench is a benchmark for offline goal-conditioned reinforcement learning, with single-task variants for standard reward-maximizing RL algorithms. We choose the single-task variants for all domains in Table 3. Following prior work (Park et al., 2024), we use the default task from each domain to tune hyperparameters. For locomotion tasks, the reward is either 1-1 or 0, where 0 indicates task completion. For manipulation tasks, the reward varies between $n_{\text{task}}$ and 0, depending on how many subtasks are complete. The datasets were collected by scripted non-Markovian policies that randomly interact with different objects in the environment. However, the successful completion of each task requires the agent to sequentially arrange those objects in a specific order (sequential reasoning), which is unseen in the dataset. Hence, the agent must be able to generalize across trajectories in the dataset and plan and sequentially combine learned manipulation skills (combinatorial generalization) (Park et al., 2025a). We use the following datasets from OGBench for each domain.
@@ E.1 Environments
In our offline-to-online experiments, we chose to conduct experiments on the following task.
  • antmaze-large-navigate-singletask-task1-v0
  • humanoidmaze-medium-navigate-singletask-task1-v0
  • cube-double-play-singletask-task2-v0
  • cube-triple-play-singletask-task1-v0
  • scene-play-singletask-task2-v0
  • puzzle-4x4-play-singletask-task4-v0
D4RL is a popular benchmark for studying offline RL. We measure the performance by the normalized returns following Fu et al. (2021). We use the following tasks from Adroit, which require dexterous manipulation within a 24-DoF action space.
@@
The cube-double and cube-triple tasks involve complex pick-and-place of colored cube blocks. The scene tasks involve long-horizon reasoning and interaction with various objects in the scene. The puzzle-3x3 and puzzle-4x4 tasks require solving the "Lights Out" puzzles using the robot arm and further test combinatorial generalization. The Adroit benchmarks involve controlling a 28-DoF hand to perform dexterous skills: spinning a pen, opening a door, relocating a ball, and using a hammer to knock a button. The antmaze and humanoidmaze tasks are designed to navigate a challenging maze by controlling an 8-DoF ant and a 21-DoF humanoid, respectively. The visual variants of these navigation tasks are challenging, as there is no low-dimensional state information, and the algorithm must directly learn from $64\times 64\times 3$ RGB images. We include challenging state-based and image-based variants for manipulation tasks as they require tackling multimodal returns and long-horizon reasoning. We also include some image-based navigation tasks to evaluate the algorithms.
E.2 Methods for Comparison
We compare Value Flows against 9 baselines, measuring the performance using success rates (OGBench) and discounted cumulative returns (D4RL). Specifically, we first compare to BC, IQL (Kostrikov et al., 2021a), and ReBRAC (Tarasov et al., 2023) as representative methods that learn scalar Q values with a Gaussian policy. The second set of baselines is state-of-the-art methods that learn scalar Q values with a flow policy (FBRAC, IFQL, FQL) (Park et al., 2025c). We also include comparisons against prior distributional RL methods that model the return distribution as a categorical distribution (C51 (Bellemare et al., 2017)) or a finite number of quantiles (IQN (Dabney et al., 2018), and CODAC (Ma et al., 2021)). We use a flow policy for all these distributional RL baselines. Following prior work (Park et al., 2025c), we report means and standard deviations over 8 random seeds for state-based tasks (4 seeds for image-based tasks). We describe the details of each method below.